### v1

# Can ChatGPT Rival Neural Machine Translation? A Comparative Study

Abstract
Inspired by the increasing interest in leveraging large language models for translation, this paper evaluates the capabilities of large language models (LLMs) represented by ChatGPT in comparison to the mainstream neural machine translation (NMT) engines in translating Chinese diplomatic texts into English. Specifically, we examine the translation quality of ChatGPT and NMT engines as measured by four automated metrics and human evaluation based on an error-typology and six analytic rubrics. Our findings show that automated metrics yield similar results for ChatGPT under different prompts and NMT systems, while human annotators tend to assign noticeably higher scores to ChatGPT when it is provided an example or contextual information about the translation task. Pairwise correlation between automated metrics and dimensions of human evaluation produces weak and non-significant results, suggesting the divergence between the two methods of translation quality assessment. These findings provide valuable insights into the potential of ChatGPT as a capable machine translator, and the influence of prompt engineering on its performance.
Can ChatGPT Rival Neural Machine Translation? A Comparative Study
Zhaokun ${}^{\dagger}$, Ziyin ${}^{\ddagger}$thanks: daenerystargaryen@sjtu.edu.cn ${}^{\dagger}$ School of Foreign Languages, Shanghai Jiao Tong University ${}^{\ddagger}$ Department of Computer Science and Engineering, Shanghai Jiao Tong University
1 Introduction
Against the backdrop of translation automation and the advances in translation technology, neural machine translation (hereafter NMT) has been widely investigated under both academic and professional scenarios (Gaspari et al., 2015). Scholars have discussed its applications across different domains, and found it to fare well in both non-literary and literary texts, demonstrating satisfying quality and even creativity to some degree (Hu and Li, 2023).
In recent years, the emergence of large language models (LLMs) has further revolutionized the relationship between translation technologies and translators, and has opened up new possibilities for interactive translation using customized prompts, making proficiency in using AI tools an integral part of digital literacy (Godwin-Jones, 2022). Among them, ChatGPT, an AI-powered LLM designed and trained by OpenAI1, has demonstrated remarkable capabilities in a wide range of tasks, including translation among others (Hendy et al., 2023; Lai et al., 2023; Hendrycks et al., 2021; Srivastava et al., 2023; Zhao et al., 2023; Zhang et al., 2023). Recent studies have shown that ChatGPT is in par with or surpasses mainstream NMT engines in translating news, user comments, conversation, social media posts, and scientific texts (Jiao et al., 2023; Hendy et al., 2023; Wang et al., 2023; Peng et al., 2023; Lu et al., 2023; Chen et al., 2023; He et al., 2023; Kocmi and Federmann, 2023; Jiang et al., 2023).
The earlier explorations focusing on translation quality assessment (TQA) of machine translation have demonstrated the promising role of ChatGPT in generating quality translations. However, most of these studies are conducted on huge translation corpora, and the scholars mostly rely on automated metrics and scarcely resort to human evaluation, which only serves as a complement to automated metrics.
Figure 1: Total error penalty (left) and proportion of error severity (middle) assigned by human annotators to different translations. Right: correlation coefficients between human evaluation and automated metrics.
As NMT systems and LLMs keep making improvements, these measures may fail to fully capture the quality of machine translation. In particular, we have limited understanding of what aspects or dimensions other than content overlap can really define high-quality machine translations and distinguish them from those low-quality ones. There is a lack of metrics that are capable of measuring translation quality as humans do, who pay attention to dimensions such as intended audience, cultural sensitivity, adherence to translation norms, textual coherence, and practicality. Also, previous research is often conducted on extensive, publicly available datasets that are crawled from the internet and encompass diverse domains [^1] In this work, the terms “domain” and “’register” are used interchangeably. The former is preferred by the AI community. The later is preferred by linguistics literature., genres, and text types. There seems to be a scarcity of high-quality, domain-specific parallel corpora that would enable more focused and specialized quality assessment in machine translation.
Thus, we focus on a specific domain, and construct our major research questions as follows:
  1. 1.
    How well do NMT systems and ChatGPT perform under automated assessment?
  2. 2.
    How well do NMT systems and ChatGPT perform under human evaluation?
  3. 3.
    To what degree do automated metrics align with human evaluation?
Essentially, RQ1 and RQ2 aim to provide a well-rounded assessment of translations from NMT systems and ChatGPT using both automated metrics and human evaluation. Based on RQ1 and RQ2, RQ3 taps into the inter-correlation between automated metrics and human scores. For RQ1 and RQ2, we find that ChatGPT excels under human evaluation and semantic-aware automatic evaluation (see Figure 1 and Table 2). For RQ3, we find that the two methods may exhibit divergences when approaching machine-translated texts, contrary to Lu and Han (2023)’s findings that automated metrics can show moderate to strong correlations with human-assigned scores in assessing interpreting outputs, possibly due to the inherent differences between interpreting and translation.
2 Related Work
2.1 Improving the Translations of LLM via Prompt Engineering
Large Language Models (LLMs) represented by OpenAI’s ChatGPT have advanced remarkably in recent years, showcasing emergent abilities such as in-context learning and Chain-of-Thought reasoning (Brown et al., 2020; Chowdhery et al., 2022; Wei et al., 2022a, b; Wang et al., 2022), and several studies have explored the influence of prompting strategies on the translation performance of LLMs (Jiao et al., 2023; Hendy et al., 2023; Wang et al., 2023; Peng et al., 2023; Lu et al., 2023; Chen et al., 2023; He et al., 2023). It has been found that with carefully written prompts, LLMs’ translation quality can gain noticeable improvement, though such better performance is not seen in every scenario (Kocmi and Federmann, 2023; Lu et al., 2023). For instance, to inject cultural awareness into LLMs such as ChatGPT, Yao et al. (2023) propose “cultural knowledge prompting”. Likewise, Mu et al. (2023) use translation memories as part of the prompts to enhance the translation performance of LLMs.
2.2 Translation Quality Assessment
Much attention has been devoted to the field of Translation Quality Assessment (TQA) in the past few decades, with both empirical and theoretical discussions increasing markedly. Particularly, many studies have emphasized the importance of MT evaluation since MT has already become an integral instrument for translation practitioners, in both translator training and translation education. Coupled with this process is the rapid development of NMT engines and latelt the emergence of chatbots represented by ChatGPT, which poses great challenges to the applicability and validity of various assessment metrics and methods.
Translation quality, however, has in fact no commonly agreed definition, as “theorists and professionals overwhelmingly agree there is no single objective way to measure quality” (Drugan, 2013). For those who share the belief that a translation should resemble the source text as closely as possible, quality is necessarily entangled with the concept of “equivalence”. However, as translation studies experienced the cultural turn, this approach faces much criticism from the descriptive approach, which holds that emphasis should be laid on observing and describing the way translations are actually carried out, and how they are influenced by socio-cultural factors. The fit-for-purpose approach advocated by functionalists stresses that the notion of quality should be decided by how translations are received by end-users in real-life contexts.
When MT becomes the subject of TQA, it can also draw on the above conceptions of quality. However, since MT is application-oriented in nature, calculability and measurability should be at the core of its quality assessment. One widely adopted view of MT quality is proposed by Koby et al. (2014), which regards accuracy and fluency as the two determinants of translation quality. Alignment with human translation is also stressed, and Hassan et al. (2018) argue that only when MT shows no significant difference from human translations as measured by some scores can it be counted as high-quality translation.
Despite the diversity of frameworks to address the evaluation of MT quality, we follow Chatzikoumi (2020) to roughly divide approaches to TQA for MT into two lines: automated metrics and human evaluation. Automated metrics are of high efficiency, speed, consistency, and cost-effectiveness, and thus are widely adopted in both the academia and the industry. Within this line, metrics that compare a given translation to one or more reference translations are reference-based. They measure the similarity between the candidate MT and the reference(s) based on various linguistic and statistical features. Examples of reference-based metrics include BLEU (Bilingual Evaluation Understudy, Papineni et al., 2002), TER (Translation Edit Rate, Snover et al., 2006), ChrF (Character n-gram F-score, Popovic, 2015), and METEOR (Metric for Evaluation of Translation with Explicit Ordering, Banerjee and Lavie, 2005).
Reference-free metrics, unlike reference-based metrics that rely on a set of high-quality reference translations, assess the quality of translations based solely on the characteristics of the translated text itself. These metrics aim to measure translation quality in a more independent and self-contained manner. They often focus on linguistic properties, statistical patterns, or other features of the translation, and are particularly useful in scenarios where high-quality reference translations are not readily available or when evaluating translations with specific requirements that do not align with existing references. Typical reference-free metrics include YiSi-2 (Lo et al., 2018), COMET-QE (Rei et al., 2021), and QuestEval (Scialom et al., 2021).
Human evaluation can also be categorized into different types based on the methodology used. In rubric scoring (Bentivogli et al., 2018), human evaluators directly assign a score for translations within a fixed rating scale according to predefined criteria such as fluency, adequacy, accuracy, grammar, style, and overall coherence. Their judgments are typically combined to provide a final quality score. Another way of human evaluation is based on ranking (Bojar et al., 2016), which involves comparing and ranking multiple translations to decide which is superior or more suitable for a given purpose or context. As for error-based evaluation, errors or issues present in translations are identified and categorized according to pre-determined error typology. Human evaluators analyze the translations and annotate specific errors, taking into consideration dimensions that include accuracy, fluency, terminology, style, and so on. The most representative error typologies are DQF (Dynamic Quality Framework) proposed by TAUS in 2011, MQM (Multidimensional Quality Metrics) developed by QTLaunchPad, and the harmonized DQF-MQM taxonomy (Popović, 2018). Finally, post-editing can also be used for evaluation purposes, since the posting-editing efforts by a human translator to render a MT “good enough” and “deliverable” implicitly reflects the quality of the raw translation (Massardo et al., 2016).
2.3 Comparive Studies of LLM Translations and NMT
As a promising contender to dedicated NMT systems, LLMs have attracted much attention from TQA-related studies that compare their performance against NMT engines using both automated metrics and human assessment. For example, Jiao et al. (2023) compare the translation performance of ChatGPT and GPT4 against three NMT systems: DeepL, Google Translate, and Tencent TranSmart using four automated metrics. They find that ChatGPT and GPT4 perform comparably well to NMT engines in specific European languages, but not in others. They also find that ChatGPT and GPT4 excel in translating spoken texts, but not in some specialized fields, such as abstracts of academic papers. Similarly, Hendy et al. (2023) show that for high-resource language pairs such as English and French, ChatGPT could exhibit state-of-the-art translation capabilities matching or even surpassing the mainstream NMT engines, while Jiang et al. (2023) extract a wide range of linguistic features of translations in political domains, and find that the translations of ChatGPT are closer to NMT than to human translations. Karpinska and Iyyer (2023) show that when translating paragraph-level texts, ChatGPT produces fewer mistranslations, grammatical errors, and stylistic inconsistencies compared to Google Translate.
The existing researches have confirmed the strong capacity of LLMs in translating high-resource languages such as English and German. However, we are not fully aware of their competence in understanding and translating middle and low-resource languages, such as Chinese. Also, most of the assessments are conducted on publicly available corpora such as WMT datasets, leaving more specialized translations - such as diplomatic discourse - as under-investigated domains awaiting more investigation (Jiang et al., 2023).
3 Methodology
3.1 Corpus
The corpus in this study consists of 6,878 pieces of Spokesperson’s Remarks on regular press conferences, which contains questions proposed by foreign reporters and answers from the Chinese spokespersons centering several foreign affairs at a range of press conferences. The corpus includes 17,837 parallel sentences in total, which amount to 642K Chinese tokens and 450K English tokens. We focus on Chinese-to-English discourse out of three considerations: (1) data availability, since all the textual materials are easily found online; (2) high quality, as the human translation is performed by professional institutional translators; (3) relative complexity, because the sources texts contain many idiomatic expressions and political terms only seen in Chinese political contexts, which may pose challenge to NMT engines and ChatGPT.
All of the questions are asked in English, and answers delivered by the spokespersons are in Chinese and then translated into English by institutional translators. Texts in our dataset are in fact transcripts of these press conferences, with some adjustments of wordings and contents, as well as corrections of speaking errors conducted by professional editors to be in line with the requirements of government websites. Since what the Chinese spokespersons express to the outside world concerns China’s national interest and stance, their English translations are carefully curated and of premiere quality. All the materials can be directly downloaded from the official website of the Ministry of Foreign Affairs of the People’s Republic of China [^2] https://www.mfa.gov.cn/eng/.
Method Prompt
0-shot
Please translate the following Chinese sentence {source} into English. You should only output the translation.
Please translate the following Chinese sentence {source} into English. You should only output the translation.
1-shot
Here is an example for you:
{source}
{reference}
Context
Please translate the following Chinese sentence {source} into English. It comes from a press conference in which a Chinese official interact with foreign reporters. The Chinese sentence is in the domain of politics. You should only output the translation.
Table 1: Prompts for querying ChatGPT.
3.2 Translation Tools
For NMT systems, we consider Microsoft Translate, Google Translate, and DeepL. After the completion of machine translation, we conducted a manual examination of the outputs to verify that they were system errors.
For LLM, we use GPT-3.5-Turbo, more commonly known as ChatGPT [^3] We use gpt-3.5-turbo-0613 API, accessed on November 2nd, 2023.. We handcraft three prompts: 0-shot, 1-shot, and a third prompt with additional information about the domain and context, as shown in Table 1.
3.3 Evaluation Methods
We adopt both automated metrics and human evaluation to balance efficiency and quality. Four widely-adopted automated metrics were selected. For human evaluation, we choose the integrated MQM-DFQ error typology and a six-dimensional analytic rubric scoring method.
3.3.1 Automated Metrics
BLEU involves a simple calculation of the number of n-gram matches between the MT output and one or multiple reference translations, with a penalty for unreasonably short translations. ChrF computes the F-score of character n-gram overlap instead. BERTScore leverages representations from pre-trained language models to compute the cosine similarity between translations and references, while COMET is similarly based on pre-trained language models, but also takes the source sentence into consideration during the scoring process.
We calculate BLEU, chrF, and BERTScore using torchmetrics [^4] https://github.com/Lightning-AI/torchmetrics. COMET is computed using Unbabel-COMET [^5] https://github.com/Unbabel/COMET.
3.3.2 Human Evaluation Based on Error Typology
Multidimensional Quality Metrics (MQM) is a hierarchical and specifications-based framework for TQA (Lommel et al., 2013) that considers different dimensions of translation quality simultaneously in a comprehensive and systematic manner. Similarly, the DQF error typology developed by TAUS further contributes to this endeavor by providing a standardized classification system for identifying and categorizing errors in translations.
The similarity between MQM and DQF makes their marriage possible. The integrated DQF-MQM typology aims to provide a unified approach to TQA. It comprises seven primary error types, several subtypes under each primary type, and 4 severity levels (critical, major, minor, neutral) to address translation issues related to accuracy, locale convention, terminology, style, design, fluency, verity, and other issues.
We choose accuracy, fluency, style, and terminology to form a subset of error typology to conduct human evaluation. To ensure the evaluation quality, annotators are provided with a detailed guideline, which can be found in the Appendix A. It covers specifications of the textual materials, error typologies used in this study, severity levels, and penalty points. Our guideline is drafted in accordance with the official file provided by TAUS [^6] https://info.taus.net/dqf-mqf-error-typology-template-download.
3.3.3 Human Evaluation Based on Analytic Rubric Scoring
Analytic rubric scoring is another method widely adopted in TQA research. It is founded on the assumption that the overall concept of quality can be broken down into individual components, and typically comprises several sub-scales addressing separate dimensions of translation (Lu and Han, 2023). To complement the error typology-based evaluation, we propose six analytic rubrics to capture translation quality from different perspectives, encompassing dimensions of (1) coherence, (2) adherence to norms, (3) style, tone, and register appropriateness, (4) cultural sensitivity, (5) clarity, and (6) practicality. We use a 7-point Likert Scale and asked annotators to assign scores to the translation outputs for each rubric.
Considering time and costs, for each origin of translation (translations by ChatGPT under three prompts and by three NMT systems), we randomly sample 50 pieces of text to conduct human evaluation. To avoid pre-conceived judgements or biases on the quality of these translation tools, we do not inform the annotators of the origin of these translation samples beforehand. We recruited 6 annotators in total, all post-graduate students majoring in translation and interpreting (MTI). Each annotator is assigned 100 pieces of texts from all the six origins, and thus each text is annotated by two annotators. In terms of the error typology-based evaluation, the two annotations yield an average Cohen’s Kappa of 0.73. Cases of disagreement (including the error type and the severity level) are settled through discussion and come to unanimous results. For the analytic rubric scoring, we average the scores assigned by the two annotators in each dimension.
Model
BLEU
Chrf
BERTScore
COMET
ChatGPT 0-shot
23.82
55.83
96.03
84.19
ChatGPT 1-shot
25.08
55.83
96.14
84.64
ChatGPT with context
23.98
56.19
96.38
85.30
Google Translate
25.28
55.86
96.12
83.02
MS Translate
29.16
59.17
96.15
84.41
DeepL
26.39
57.39
96.18
83.92
Table 2: Automated metrics of ChatGPT and NMT systems.
3.4 Analysis of Scores from Automated Metrics and Human Annotators
To answer RQ1, we compute the average scores of the four automated metrics. Their descriptive statistics can be found in Appendix B. To answer RQ2, we first provide annotation results guided by the MQM-DQF error typology, including the average error penalties, the frequencies of error types and subtypes, and the distribution of error severities. Then we show the results of analytic rubric scoring on six dimensions. The detailed descriptive statistics of human evaluation are listed in Appendix C. To answer RQ3, we calculate the Pearson’s correlation coefficients at the significance level of 0.05 to examine the inter-relationship between automated metrics and human scores.
4 Results and Analysis
This section provides results of automated metrics and human evaluation. Section 4.1 shows the performance of ChatGPT and NMT engines measured by four automated metrics. The first part of Section 4.2 centers on the results of human evaluation, displaying error penalties, the distribution of error severities and error types in different translation outputs. The second part displays scores assigned by human annotators in six analytic rubrics. Section 4.3 shows the inter-correlation between automated metrics and human evaluation.
4.1 Automated Metrics
The results of ChatGPT and NMT systems’ translations are presented in Table 2. Compared with the 0-shot scenario, providing ChatGPT with an example or contextual information only brings slight increase in BLEU score. There is no noticeable change on other evaluative metrics, indicating that ChatGPT under the 0-shot condition has already demonstrated strong capability in handling this specific translation task. The high BERTScore and COMET scores prove that in terms of semantic accuracy, ChatGPT has no difficulty in understanding the characteristic expressions in the domain of diplomatic discourse, and can deliver relatively faithful English translations even under the 0-shot condition. Among the three NMT systems, MS Translate gains the highest scores on 3 out of 4 metrics, only slightly lagging behind DeepL on BERTScore. However, the differences among these systems are not marked.
ChatGPT Google MS DeepL
0-shot 1-shot w. context
Accuracy Addition 1 0 1 0 0 0
Omission 1 0 2 3 0 0
Mistranslation 4 7 4 6 6 10
Over-translation 1 1 1 2 2 2
Under-translation 5 1 2 0 3 4
Total 11 8 9 12 11 15
Fluency Punctuation 0 0 1 2 0 0
Grammar 6 5 8 5 8 8
Total 6 5 8 7 8 8
Terminology Wrong terms 9 6 6 7 4 7
Inconsistent usage 0 0 0 0 0 0
Total 9 7 7 7 5 8
Style Inconsistent style 0 1 0 1 1 1
Awkward 16 2 3 2 3 2
Unidiomatic 3 13 7 9 4 9
Total 19 17 12 13 9 13
Other Errors 0 1 0 1 0 0
Table 3: Major types and subtypes of errors in different translations.
Figure 2: Human-assigned scores for each analytic rubric.
4.2 Human Assessment
The total error penalty assigned by human annotators to each translation system is plotted in the left subfigure of Figure 1. We observe that ChatGPT under the 1-shot condition is assigned the lowest penalties, indicating its highest translation capacity among the translation outputs listed, while ChatGPT under the 0-shot condition performs much worse. This suggests that providing even a small amount of training data, in this case only one example, significantly improves the translation capability of ChatGPT. Error penalty in translations by ChatGPT when it is given contextual information is the second lowest, showing that incorporating context, possibly through the use of context-aware prompts, can improve the translation quality provided by ChatGPT. On the other hand, Google Translate, MS Translate, and DeepL have comparably high error penalties, indicating similar translation performance among these widely used machine translation systems.
The middle subfigure of Figure 1 displays the proportion of errors in each level of severity. Among NMT engines, DeepL contains the highest percentage of neutral errors and least percentage of major and critical errors, indicating that it demonstrates better capability compared with the other two NMT engines. In contrast, Google Translate sees the highest proportion of major errors in its translations. As for ChatGPT, most of its errors under the 0-shot condition are minor and major errors. Giving it an example significantly decreases the proportion of its critical errors and increases its neutral errors. Providing it with contextual information yields a similar effect.
Table 3 shows the occurrence of errors under each major and sub-category. Overall, stylistic errors occur most frequently in most translations, followed by errors related to accuracy. While most errors of ChatGPT are attributed to style, accuracy poses the most challenge to NMT systems. Particularly, translations by ChatGPT are identified significantly more awkward and unidiomatic expressions compared with NMT systems.
In terms of analytic rubrics, Figure 2 shows that ChatGPT under the 1-shot setting consistently receives the highest scores across all the rubrics, indicating its overall superiority under human evaluation. In contrast, ChatGPT in the 0-shot scenario ranks last in five out of six dimensions (adherence to norms, practicality, clarity, cultural sensitivity, as well as style, tone, and register appropriateness), which suggests that providing it with an example containing high-quality human translation contributes to boosting its translation performance markedly. Exposing ChatGPT to contextual information also serves to improve its translation quality in most cases (four out of six dimensions). The three NMT systems receive similar scores, showing that human annotators do not observe notable differences among their translation outputs.
4.3 Correlation between Automated Metrics and Human Assessment
Based on the results in Section 4.1 and ref 4.2, we calculate the correlation coefficients between human-assigned scores and scores from automated metrics. The results are shown in the right subfigure of Figure 1. We observe that in general, BERTScores align with human-assigned scores most closely, showing an average correlation coefficient of 0.16. BLEU scores deviate most from human evaluation, with a mere 0.10 correlation coefficient on average. However, the majority of these correlations are not statistically significant (the p values can be found in Appendix D), suggesting that human evaluation and automated metrics tend to focus on different aspects of translation quality. In terms of aspects of human evaluation, error penalty shows the strongest correlation with automated scores in general, while cultural sensitivity demonstrates the weakest correlation.
5 Discussion
From the results of automatic assessment, we can see that when measured by BLEU and Chrf, ChatGPT is overshadowed by NMT systems. This suggests that translations generated by ChatGPT tend to deviate more from the references in terms of the n-gram matches than NMT system, and ChatGPT may struggle to produce translations that precisely replicate the wording and phrasing of the references. However, its performance shines when evaluated using BERTScore and COMET score. These metrics take into account the semantic similarity and fluency of translations, rather than focusing solely on n-gram overlap. ChatGPT’s ability to challenge or even surpass NMT systems when measured by these two metrics suggests that despite the deviations from the reference translations, its translations exhibit a strong semantic closeness to the references.
This observation aligns with previous researches on generative LLMs, which have shown that these models excel at capturing semantic relationships and generating coherent text. The underlying architecture of ChatGPT grants it exceptional context awareness and ability of language understanding, which enables its to produce translations that are more semantically aligned with the references, even if they may differ in specific word choices or phrase structure. It is worth noting that although BLEU and Chrf scores are widely adopted for evaluating translation quality, they have limitations, particularly when applied to language models like ChatGPT, which prioritize semantic coherence over exact phrasing. On the other hand, BERTScore and COMET provide a more nuanced evaluation by considering semantic aspects and fluency. These metrics are better at reflecting the strengths of ChatGPT in translating context-dependent texts.
Another finding is that different prompting strategies do not exert significant boost to ChatGPT’s translation performance, which can be attributed to two possible reasons. First, ChatGPT in the 0-shot scenario can already understand the semantic meaning of the original text perfectly, so that providing it with either extra information will not significantly affect its translation capability. Second, these automated metrics fail to capture the improvement brought by customized prompts. Drawing from results of human evaluation, the second explanation holds more feasibility, which will be further discussed later.
In terms of human evaluation, we observe from the overall error penalties (see Figure 1) that ChatGPT tends to demonstrate strong plasticity in its translation performance: providing it with only one example or relevant contextual information can greatly reduce error penalties in its translation outputs and lead to a significant improvement in translation quality.
This is further corroborated by evidence from scores of six analytic rubrics (see Figure 2), which demonstrates the variability of the performance of ChatGPT under different conditions. When exposed to a single example, ChatGPT consistently receives the highest scores across all the analytic rubrics, indicating its superior performance as evaluated by human annotators. In contrast, ChatGPT without any example (0-shot condition) ranks last in five out of six dimensions: adherence to norms, practicality, clarity, cultural sensitivity, and style, tone, and register appropriateness. This indicates that ChatGPT without the guidance of an example struggles in these areas and relies on the presence of a reference translation to enhance its performance. Similarly, exposing ChatGPT to contextual information improves its translation quality across all the six rubrics, showing that the inclusion of context aids ChatGPT in generating more contextually appropriate translations. This informs us the importance of crafting appropriate and relevant prompts to fully uncover the potential of ChatGPT in generating high-quality translations. On the other hand, the three NMT systems receive similar scores across the assessed dimensions, implying that human annotators do not observe significant differences in translation quality among these NMT systems.
In addition, NMT engines are more prone to mistakes compared with ChatGPT. This corresponds to previous findings reported by Manakhimova et al. (2023), who investigate the ability of GPT4 and NMT systems in handling challenging translation issues. Their study shows that GPT4 outperforms NMT engines in most cases, displaying fewer mistakes and higher accuracy. This is further supported by the distribution of errors (see Table 3), which shows that accuracy poses the greatest difficulty for NMT systems in general.
In contrast, ChatGPT exhibits style errors as its primary challenge. In particular, translations by ChatGPT are noted for containing more awkward and idiomatic expressions compared to NMT systems, but providing relevant context notably reduced its style-related errors. ChatGPT’s style errors may stem from its language generation capabilities, as it tends to prioritize semantic coherence over precise phrasing, leading to stylistic variations. The awkward and idiomatic expressions may arise from its tendency to generate creative and diverse outputs, which can sometimes result in less conventional or less fluent translations.
Finally, the Pearson’s correlations between human-assigned scores and automated metrics are weak and non-significant, though some specific dimensions show slightly higher correlations than others. This is somehow contradictory to findings in Lu and Han (2023), possibly because assessing interpreting is different from assessing translation, with the latter having higher requirements on aspects other than accuracy. Our results highlight the challenges in relying solely on automated scores to assess translation quality, as they may not fully capture the intricacies of dimensions such as adherence to norms, cultural sensitivity, clarity, and practicality. Human evaluation as a more contextual and culture-specific way of assessment is necessary to obtain a comprehensive understanding of translation quality.
What implications do these results have for future research in TQA? First, we should acknowledge the limitations of traditional metrics. Our findings highlight that metrics like BLEU and chrF, which primarily focus on n-gram overlap and exact phrasing, cannot effectively capture the superior performance of language models such as ChatGPT in a range of critical dimensions for high-quality human-like translations. This calls for the development and adoption of more nuanced evaluation metrics that consider cultural aspects and contextual appropriateness. Next, carefully crafted and customized prompts are needed to fully unleash the great potential of ChatGPT as a capable machine translator. Our results show that providing a single example or some relevant contextual information can greatly reduce its errors and drive up its scores in all the analytic rubrics. This highlights the practical significance of tailoring prompts to guide the generation process and enhance the translation quality of ChatGPT.
6 Conclusion
This study compares the translation quality of ChatGPT and three NMT engines by using both automated metrics and human evaluation. For the former, we compute four widely-adopted metrics - BLEU, chrF, COMET, and BERTScore - and find that they fail to distinguish high-quality and lower-quality translations. For the latter, we conduct annotation based on both the integrated MQM-DQF error typology and six analytic rubrics. Results show that exposing ChatGPT to one example or context-relevant information greatly boosts its performance under all dimensions of human evaluation. To examine the correlation between automated metrics and human evaluation, we calculate the pairwise Pearson’s correlation coefficients. The weak and non-significant results overall demonstrate that human understanding of translation quality is significantly different from what is captured by automated metrics.
In light of the findings discussed above, we suggest two directions to advance researches in TQA: Firstly, further exploration is needed to develop more effective evaluation metrics that can better capture the nuances of translation quality, particularly in terms of coherence, clarity, practicality adherence to norms, cultural sensitivity, as well as appropriateness of style, tone, and register. Secondly, it is crucial to continue investigating the role of proper prompts and contextual information in improving the performance of language models like ChatGPT, which exhibit strong context awareness and language understanding capabilities.
Appendix A Guidelines for Human Assessment under the MQM Error Typology
A.1 General Instructions
You will be assessing translations at the sentence level. Each translated sentence is aligned with its corresponding source text. You have the flexibility to revise previous annotations as needed. There are two tasks for you to finish. The first one is error-analysis-based. Your task is to identify errors within each translated sentence, with a maximum limit of five errors. If there are more than five errors, focus on marking the five most severe ones. In cases where the translation is severely distorted or unrelated to the source, mark a single Non-translation error that covers the entire segment. To identify an error, highlight the relevant portion of the translation and choose a category/sub-category and severity level from the available options. When identifying errors, please be as fine-grained as possible. For instance, if a sentence contains two mistranslated words, record them as separate mistranslation errors. If a single section of text has multiple errors, indicate the most severe one. Please pay particular attention to the context when annotating. If a translation may be questionable on its own but fits within the context, it should not be considered erroneous. Conversely, if a translation might be acceptable in some contexts, but not for the current sentence, mark it as incorrect. There is a special error category called Non-translation, which can only be used once per sentence and should encompass the entire sentence. If Non-translation is selected, no other errors should be identified. The second task is impression-based. You need to report your level of agreement with 7 statements on a 7-point Likert scale, in which 1 means “completely disagree”, and 7 means “completely agree”. The statements are the following: 1. Coherence: The translation flow is mostly smooth and coherent. There is no logical disconnection or meaning inconsistency. 2. Adherence to norms: The translation fulfills the common standards, requirements, and norms of political translation. 3. Style, tone, and register appropriateness: The translation is consistent in style, tone, and register with the source text. For example, if the source text has a formal tone and sophisticated style, the translation also reflects that formality and sophistication. 4. Cultural sensitivity: The translation demonstrates cultural sensitivity. It suitably conveys culture specific items (CSI), humor, and other cultural nuances in a way that is understandable and relatable to the target audience. 5. Clarity: The translation is clear and easily understandable to the target audience. It does not contain ambiguities, excessive jargon, or overly complex language that may hinder comprehension. 6. Practicality: The translation can be directly put for actual use. In this case, it can be put on the government website for people across the world to read.
A.2 Specifications
We select a portion of parameters from the ASTM F2575-14 specifications [^7] https://www.astm.org/f2575-14.html to describe what is expected of the translation.
Source-content information
Source language: Chinese Text type: political and diplomatic texts remarks from the Chinese spokesmen. Audience: Chinese readers who intend to know China’s stance on a range of important foreign affairs (e.g., reporters, politicians, and diplomats). Purpose: to deliver China’s stance and attitudes on a range of important foreign affairs. Volume: you will be given 100 sentences. Complexity: usually written in a relatively complex and formal style, commonly found in official statements or diplomatic contexts. Origin: official website of the Ministry of Foreign Affairs of The People’s Republic of China (https://www.fmprc.gov.cn/fyrbt_673021/dhdw_673027/index_1.shtml)
Target content requirements
Target language: English Audience: international readers who intend to know China’s stance on a range of important foreign affairs (e.g., reporters, politicians, and diplomats). Purpose: to deliver China’s stance and attitudes on a range of important foreign affairs. Format: written texts displayed on the government website, which are transcribed and carefully edited from spoken remarks. Style: in a formal, official, and often assertive tone commonly found in official statements or diplomatic discourse.
A.3 Error Typology, Severity Levels and Penalty Levels
Error Typology
Error Type Subtype Definition Accuracy The target text does not accurately reflect the source text, allowing for any differences authorized by specifications. Addition The target text includes text not present in the source. Omission Content is missing from the translation that is present in the source. Mistranslation The target content does not accurately represent the source content. Over-translation The target text is more specific than the source text. Under-translation The target text is less specific than the source text. Fluency Issues related to the form or content of a text, irrespective as to whether it is a translation or not. Punctuation is used incorrectly (for the locale or style). Spelling Issues related to spelling of words. Grammar Issues related to the grammar or syntax of the text, other than spelling and orthography. Inconsistency The text shows internal inconsistency. Link/cross-reference Links are inconsistent in the text. Terminology A term (domain-specific word) is translated with a term other than the one expected for the domain or otherwise specified. Wrong terms Use of term that it is not the term a domain expert would use or because it gives rise to a conceptual mismatch. Inconsistent use of terminology Terminology is used in an inconsistent manner within the text. Style The text has stylistic problems. Inconsistent style Style is inconsistent within a text. Awkward A text is written with an awkward style. Unidiomatic The content is grammatical, but not idiomatic. Other Any other issues.
Severity Levels
Non-translation The translation is severely distorted or unrelated to the source. Critical Errors that may carry health, safety, legal or financial implications, violate geopolitical usage guidelines, damage the entities’ reputation, or which could be seen as offensive. Major Errors that may confuse or mislead the readers due to significant change in meaning or because errors appear in a visible or important part of the content. Minor Errors that don’t lead to loss of meaning and wouldn’t confuse or mislead the readers but would be noticed, would decrease stylistic quality, fluency or clarity, or would make the content less appealing. Neutral Used to log additional information, problems or changes to be made that don’t count as errors, e.g., they reflect a reviewer’s choice or preferred style, they are repeated errors or instruction/glossary changes not yet implemented, a change to be made that the translator is not aware of.
Penalty Levels
Non-translation 100 Deduct the penalty points for non-translation errors
Critical errors 25 Deduct the penalty points for critical errors
Major errors 10 Deduct the penalty points for major errors
Minor errors 1 Deduct the penalty points for minor errors
Neutral errors 0 Deduct the penalty points for neutral errors
Appendix B Descriptive Statistics of Scores Using Automated Metrics
The descriptive statistics of automated metrics (BLEU, chrF, COMET, and BERTScore) are given in Table 4, 5, 6, and 7 respectively.
Mean Std. Min Max Kurtosis Skewness
ChatGPT 0-shot 0.228 0.196 0.000 1.000 0.211 0.605
ChatGPT 1-shot 0.251 0.213 0.000 1.000 0.317 0.662
ChatGPT w. context 0.235 0.200 0.000 1.000 0.172 0.607
Google Translate 0.243 0.209 0.000 1.000 0.306 0.654
MS Translate 0.264 0.217 0.000 1.000 -0.089 0.528
DeepL 0.292 0.236 0.000 1.000 -0.039 0.562
Table 4: BLEU statistics.
Mean Std. Min Max Kurtosis Skewness
ChatGPT 0-shot 0.558 0.152 0.076 1.000 0.214 -0.333
ChatGPT 1-shot 0.586 0.092 0.173 1.000 0.357 -0.178
ChatGPT w. context 0.563 0.154 0.067 1.000 0.215 -0.314
Google Translate 0.559 0.167 0.000 1.000 0.754 -0.421
MS Translate 0.592 0.169 0.065 1.000 0.039 -0.147
DeepL 0.574 0.163 0.065 1.000 0.077 -0.282
Table 5: ChrF statistics.
Mean Std. Min Max Kurtosis Skewness
ChatGPT 0-shot 0.842 0.059 0.441 0.986 5.068 -1.512
ChatGPT 1-shot 0.846 0.061 0.441 0.986 3.902 -1.296
ChatGPT w. context 0.841 0.062 0.368 0.986 6.591 -1.714
Google Translate 0.830 0.084 0.269 0.986 13.968 -2.824
MS Translate 0.844 0.068 0.341 0.986 4.771 -1.386
DeepL 0.839 0.064 0.441 0.986 3.549 -1.166
Table 6: COMET statistics.
Mean Std. Min Max Kurtosis Skewness
ChatGPT 0-shot 0.960 0.007 0.938 0.992 0.394 0.302
ChatGPT 1-shot 0.961 0.008 0.940 0.999 0.992 0.579
ChatGPT w. context 0.961 0.008 0.938 0.995 0.470 0.471
Google Translate 0.961 0.008 0.939 0.995 0.313 0.328
MS Translate 0.962 0.008 0.938 0.996 0.832 0.540
Tengxun Translate 0.962 0.008 0.939 0.995 0.283 0.408
Table 7: BERTScore statistics.
Appendix C Descriptive Statistics of Scores Assigned by Human Annotators
The descriptive statistics of annotators’ assigned scores to the six systems are given in Table 8-13.
Error penalty Coherence Adherence to norms Style, tone, and register appropriateness Cultural sensitivity Clarity Practicality Mean -4.110 4.830 4.890 5.250 4.850 4.970 4.340 Std. -6.302 1.640 1.406 1.351 1.591 1.611 1.713 Max 10.000 2.000 2.000 2.000 1.000 1.000 1.000 Min -35.000 7.000 7.000 7.000 7.000 7.000 7.000 Kurtosis -4.373 -0.973 -1.028 -0.603 -0.917 -0.728 -1.066 Skewness -1.225 -0.444 -0.023 -0.560 -0.357 -0.563 -0.101
Table 8: Human annotation statistics of ChatGPT 0-shot.
Error penalty Coherence Adherence to norms Style, tone, and register appropriateness Cultural sensitivity Clarity Practicality Mean -3.808 5.380 5.340 5.680 5.470 5.420 4.717 Std. -7.675 1.285 1.121 1.024 1.259 1.365 1.385 Max 10.000 1.000 2.000 2.000 2.000 2.000 1.000 Min -25.000 7.000 7.000 7.000 7.000 7.000 7.000 Kurtosis -1.268 0.708 -0.138 0.297 -0.233 -0.413 -0.303 Skewness -1.218 -0.937 -0.438 -0.578 -0.604 -0.598 -0.386
Table 9: Human annotation statistics of ChatGPT 1-shot.
Error penalty Coherence Adherence to norms Style, tone, and register appropriateness Cultural sensitivity Clarity Practicality Mean -4.110 4.830 4.890 5.250 4.850 4.970 4.340 Sd. -6.302 1.640 1.406 1.351 1.591 1.611 1.713 Max 10.000 2.000 2.000 2.000 1.000 1.000 1.000 Min -35.000 7.000 7.000 7.000 7.000 7.000 7.000 Kurtosis -4.373 -0.973 -1.028 -0.603 -0.917 -0.728 -1.066 Skewness -1.225 -0.444 -0.023 -0.560 -0.357 -0.563 -0.101
Table 10: Human annotation statistics of ChatGPT with context.
Error penalty Coherence Adherence to norms Style, tone, and register appropriateness Cultural sensitivity Clarity Practicality Mean -5.660 5.260 5.220 5.290 5.170 5.110 4.384 Std. -7.907 1.474 1.345 1.328 1.464 1.645 1.658 Max 5.000 1.000 2.000 3.000 2.000 1.000 1.000 Min -25.000 7.000 7.000 7.000 7.000 7.000 7.000 Kurtosis -0.210 0.215 -0.980 -1.116 -1.221 -0.754 -0.890 Skewness -1.066 -0.911 -0.231 -0.179 -0.277 -0.588 -0.058
Table 11: Human annotation statistics of Google Translate.
Error penalty Coherence Adherence to norms Style, tone, and register appropriateness Cultural sensitivity Clarity Practicality Mean -5.540 4.788 4.717 5.152 4.889 4.768 4.384 Std. -9.440 1.728 1.597 1.466 1.684 1.778 1.748 Max 10.000 1.000 1.000 1.000 1.000 1.000 1.000 Min -35.000 7.000 7.000 7.000 7.000 7.000 7.000 Kurtosis -1.262 -0.959 -0.502 0.210 -0.554 -0.868 -1.055 Skewness -1.407 -0.397 -0.452 -0.752 -0.559 -0.489 -0.358
Table 12: Human annotation statistics of Microsoft Translate.
Error penalty Coherence Adherence to norms Style, tone, and register appropriateness Cultural sensitivity Clarity Practicality Mean -4.900 4.950 4.677 5.121 4.790 4.830 4.300 Std. -6.870 1.438 1.268 1.118 1.266 1.443 1.501 Max 0.000 2.000 2.000 2.000 2.000 1.000 1.000 Min -25.000 7.000 7.000 7.000 7.000 7.000 7.000 Kurtosis -1.411 -0.276 -0.311 -0.191 -0.283 -0.216 -0.554 Skewness -1.477 -0.547 -0.160 -0.196 -0.081 -0.410 -0.395
Table 13: Human annotation statistics of DeepL.
Appendix D Average Correlation Coefficients between Human Evaluation and Automated Metrics
See Table 14.
BLEU chrF COMET BERTScore
Error penalty 0.141 $*{}^{*}$ 0.170 $*{}^{*}$
Coherence 0.103 0.103 0.088 0.171
Adherence to norms 0.109 0.137 0.106 $*{}^{*}$
Style, tone, and register appropriateness 0.121 0.179 0.148 0.192
Cultural sensitivity 0.055 0.089 0.085 0.104
Clarity 0.085 0.134 0.095 0.150
Practicality 0.099 0.143 0.141 0.110
Average 0.102 0.133 0.119 0.159
Table 14: Correlation coefficients between human evaluation and automated metrics. $*{}^{*}$: $p<0.005$, significant in all six systems.